# Remote Sensing Spatio-Temporal Vision-Language Models: A Comprehensive Survey

### v2

@@ Abstract
The interpretation of multi-temporal remote sensing imagery is critical for monitoring Earth’s dynamic processes—yet previous change detection methods, which produce binary or semantic masks, fall short of providing human-readable insights into changes. Recent advances in Vision–Language Models (VLMs) have opened a new frontier by fusing visual and linguistic modalities, enabling spatio-temporal vision-language understanding: models that not only capture spatial and temporal dependencies to recognize changes but also provide a richer interactive semantic analysis of temporal images (e.g., generate descriptive captions and answer natural-language queries). In this survey, we present the first comprehensive review of RS-STVLMs. The survey covers the evolution of models from early task-specific models to recent general foundation models that leverage powerful large language models. We discuss progress in representative tasks, such as change captioning, change question answering, and change grounding. Moreover, we systematically dissect the fundamental components and key technologies underlying these models, and review the datasets and evaluation metrics that have driven the field. By synthesizing task-level insights with a deep dive into shared architectural patterns, we aim to illuminate current achievements and chart promising directions for future research in spatio-temporal vision-language understanding for remote sensing. We will keep tracing related works at https://github.com/Chen-Yang-Liu/Awesome-RS-SpatioTemporal-VLMs
@@
Remote Sensing, Spatio-Temporal Understanding, Vision-Language Model, Foundation Model, and Large Language Model.
@@ I Introduction
Early efforts in remote sensing image interpretation primarily concentrated on single-image analysis tasks, such as land classification [9, 10], object detection [11, 12], and semantic segmentation [13, 14]. However, single-temporal imagery reflects only the spatial surface conditions at a specific moment and fails to capture dynamic change processes like urban expansion, deforestation, or disaster aftermath. As sensing platforms and revisit frequencies have improved [15, 16, 17, 18, 19], multi-temporal imagery has become available, opening the door to remote sensing spatio-temporal image interpretation: models must understand spatial patterns (e.g., object class, spatial context) and capture temporal evolution (e.g., appearance, disappearance, growth) across time.
@@
Early research on the spatio-temporal image interpretation focused on binary change detection, which produces pixel-level masks to localize changed object regions based on temporal image sequences, such as vegetation cover changes or the emergence of buildings [20, 21]. While effective at localizing where change occurs, these approaches lack semantic insight into what has changed. Semantic change detection extended this paradigm by assigning change object labels to changed areas [22, 23, 24], yet remained confined to visual-level assessments. They struggle to achieve a higher-level semantic understanding of changes and interactive queries in a human-readable manner, such as the evolving state of objects, color of changed objects and their position relationship [25, 26, 27].
@@
In pursuit of richer, more flexible semantic understanding, some researchers have turned to Remote Sensing Spatio-Temporal Vision-Language Models (RS-STVLMs), which integrate high-level natural language modality with temporal image sequences. They develop multiple tasks such as change captioning [26, 25] and change question answering [28, 29], thereby enriching the toolkit for remote sensing spatio-temporal image interpretation. Language, as a vehicle for human communication and knowledge [30], enhances the higher-level understanding ability of RS-STVLMs. By combining language with temporal images, RS-STVLMs can not only identify spatio-temporal changes but also generate detailed descriptions and answers, facilitating multimodal interactions that extend beyond mere visual change detection.
@@
Figure 1 illustrates the timeline of some representative RS-STVLMs along with their publication dates, highlighting an active research trend since 2021. Despite this momentum, existing research remains fragmented, typically concentrating on isolated, task-specific methods, making it challenging for researchers to obtain a holistic view of the field’s progress and emerging trends. There is a pressing need for a holistic and systematic review of RS-STVLMs.
In this survey, we present the first comprehensive review of RS-STVLMs. Fig. 2 shows the structure of this survey. The survey covers the evolution of models from early task-specific models to recent general foundation models that leverage powerful large language models. We discuss progress in representative tasks, such as change captioning, change question answering, and change grounding. Moreover, we systematically dissect the fundamental components and key technologies underlying these models, and review the datasets and evaluation metrics that have driven the field. By synthesizing task-level insights with a deep dive into shared architectural patterns, we aim to illuminate current achievements and chart promising directions for future research in spatio-temporal vision-language understanding for remote sensing.
@@
Figure 1: Timeline of RS-STVLMs, illustrating the field’s rapid ascent. We outline the evolution of representative tasks and show advanced progress of how Large Language Models (LLMs) are reshaping the RS-STVLM landscape. For additional resources and daily updates, visit our GitHub page at https://github.com/Chen-Yang-Liu/Awesome-RS-Temporal-VLM
@@
Figure 3: Three key concepts in the remote sensing spatio-temporal image interpretation: (a) Binary change detection, locating areas of change without providing deeper contextual information. (b) Semantic change detection, offering insights into the types of changed objects. (c) Spatio-Temporal Vision-Language Understanding, incorporating higher-level flexible language rather than being limited to visual-level interpretation.
II Evolution of Spatio-Temporal Vision-Language Models
II-A From Visual Change Detection to Spatio-Temporal Vision-Language Understanding
@@ II-A1 Remote Sensing Change Detection
Remote sensing change detection (CD) is a fundamental spatio-temporal image interpretation task aimed at identifying differences between images acquired at different time points [31, 32, 33]. The evolution of CD methods can be viewed from both a task and a technical perspective. In the early stages, the focus was on binary change detection—locating areas of change at the pixel level without providing semantic insights into the specific types or attributes of the changes. As the field progressed, semantic change detection emerged. It extends binary methods by labeling each changed region with a specific land-cover category, thereby providing a higher level of semantic interpretation.
On the perspective of technological evolution, early CD approaches relied on traditional techniques, including algebra-based methods such as Change Vector Analysis (CVA)[34, 35], transformation-based methods like Principal Component Analysis (PCA)[36, 37] and Multivariate Alteration Detection (MAD)[38, 39], as well as classification-based approaches such as post-classification comparisons[40, 41]. These conventional methods laid the groundwork for change detection but were limited by their reliance on hand-crafted features and simple statistical models.
@@
CNN-based Methods have been the cornerstone of deep learning approaches for change detection in remote sensing [42, 43, 44]. These models leverage hierarchical convolutional operations to extract spatial change features from bi-temporal images, making them effective at capturing local textures and patterns for refined pixel-level mask prediction. Daudt et al. [42] first introduced FCNs into the CD task and presented Siamese architectures to adapt bi-temporal images. To improve change detection accuracy, many CNN-based works have been developed, such as proposing multi-scale feature fusion strategies [45, 46, 47], and designing attention modules [48, 49].
@@
Transformer-based Methods represent a significant leap forward in change detection [50, 51, 52, 53, 54]. Different from CNNs with limited receptive fields, through the self-attention mechanism, the Transformer captures long-range dependencies and contextual relationships across the entire image. As a pioneering approach, Chen et al. [50] represent the multi-temporal images as a few semantic tokens and use a Transformer to refine the visual features, revealing the change of interest. Subsequently, many methods have emerged, such as using pure Transformer [55, 56], combining transformers and CNNs to extract rich global context and local detailed information [57, 58], and lightweight Transformer [53, 59, 60]
@@
Mamba-based methods represent a recent advantage in change detection [61, 62, 63, 64, 65]. Unlike Transformer, which typically exhibits quadratic complexity to the input sequence length, Mamba achieves linear complexity while maintaining a global receptive field. ChangeMamba [62] first introduced Mamba architecture for the CD task and proposed three spatial-temporal relationship modeling mechanisms. CDMamaba [63] integrates local clues in the Mamba block and dynamically performs the interaction of bi-temporal features guided by global-local features to achieve fine-grained change detection. Zhao et al. [61] presented an Omnidirectional Selective Scan Module to selectively scan high-resolution images in multiple directions.
@@
In summary, the evolution from binary change detection to semantic change detection reflects a growing demand for a more comprehensive understanding of spatio-temporal dynamics in remote sensing imagery. Technically, the field has transitioned from traditional methods to deep learning approaches—progressing from CNN-based methods to Transformer-based models, and now to emerging frameworks like Mamba. These advancements have significantly improved the capacity to detect not only where changes occur, but also what types of changes occur, thereby laying the foundation for subsequent multimodal developments in remote sensing spatio-temporal vision-language models.
II-A2 Remote Sensing Spatio-Temporal Vision-Language Tasks
@@
Building on the advances in remote sensing change detection, recent research has expanded towards integrating both visual and linguistic modalities, forming a set of comprehensive tasks collectively termed Temporal Vision-Language Understanding tasks. These tasks not only involve detecting and localizing changes in temporal images but also involve generating natural language descriptions, interactive question answering. This evolution reflects the community’s pursuit of a deeper semantic understanding of dynamic changes in remote sensing spatio-temporal image analysis. In this section, we outline the primary tasks in RS-STVLMs, providing formal definitions.
@@
Formally, let $I=\{I_{t_{1}},I_{t_{2}},...,I_{t_{n}}\}$ represent a sequence of remote sensing images acquired at different time steps, and let $Q$ denote a textual query, prompt, or instruction. The goal of spatio-temporal vision-language models is to infer a task-specific output, which varies across different tasks.
@@
Others: A few pioneering works have explored novel spatio-temporal vision-language tasks. Text-driven Temporal Image Generation aims to generate a sequence of remote sensing images that reflects dynamic changes described in a user-provided textual prompt, such as [66, 67]. Irvin et al. [68] introduce Temporal Referring Expression task, which requires the model to identify when (in which image) a specific change described by the user occurs.
These tasks underscore the transition from early change detection methods—focused solely on visual cues—to a multi-modal approach that leverages language for deep semantic understanding. By incorporating natural language, RS-STVLMs can provide richer, more human-centered insights into dynamic Earth observation.
II-B Remote Sensing Spatio-Temporal Vision-Language Models
In this section, we introduce the methodologies underlying each of the above spatio-temporal vision-language understanding tasks, highlighting key architectures.
Figure 4: The general framework for some representative remote sensing spatio-temporal vision-language tasks: (a) Remote Sensing Change Captioning, (b) Multi-task learning of Change Detection and Captioning, (c) Remote Sensing Change Question Answering, (d) Remote Sensing Text-driven Temporal Images Retrieval, and (e) Remote Sensing Change Grounding.
@@ II-B Remote Sensing Spatio-Temporal Vision-Language Models
CNN-RNN [69] 2021.10 VGG-16 RNN First exploration of the task and experiments on a private dataset.
CC-RNN/SVM [70] 2022.08 VGG-16 RNN, SVM Published two small datasets and explored two bi-temporal feature fusion strategies.
RSICCformer [26] 2022.11 ResNet-101 Tr-Dec Proposed a large dataset LEVIR-CC and a dual-branch Transformer-based algorithm for change captioning. [Link]
PSNet [71] 2023.07 ViT-B/32 Tr-Dec Proposed multiple progressive difference-aware layers for multi-scale feature extraction. [Link]
PromptCC [72] 2023.10 ViT-B/32 GPT-2 Decoupled change caption into binary image change classification and fine-grained change perception. Proposed a multi-prompt learning strategy which can prompt the LLM to generate plausible captions. [Link]
Chg2Cap [25] 2023.11 ResNet-101 Tr-Dec Proposed attentional encoder to accurately localize the variations between diachronic image features. [Link]
ICT-Net [73] 2023.11 ResNet-101 Tr-Dec Using ResNet for multi-scale feature extraction, Interactive Change-Aware Encoder for change representation. [Link]
SITS-CC [74] 2024.03 ResNet-101 Tr-Dec Explored a new task of image time series change caption. [Link]
RSCaMa [75] 2024.05 ViT-B/32 Mamba, Tr-Dec, GPT-2 Proposed a spatio-temporal model for spatial and temporal SSM to improve temporal modeling. It introduced Mamba to this task. [Link]
SparseFocus [60] 2024.05 ResNet-101 Tr-Dec Proposed a Sparse Focus Transformer (SFT) to focus changing regions and reduce computational complexity. [Link]
SEN [76] 2024.05 ResNet with 6-channel Tr-Dec Proposed a single-stream extractor network (SEN) with contrastive pre-training to improve vision encoder. It has a lower computational cost than the dual-stream extractor. [Link]
Diffusion-RSCC [77] 2024.05 ResNet-101 Diffusion Introduced Diffusion model into this task that can learn the cross-modal data distribution between the image pairs and change captions. [Link]
@@
ChangeRetCap [79] 2024.06 ResNet-101 Tr-Dec Proposed a foundation model considering both bi-temporal captioning and bi-temporal text-driven temporal images retrieval tasks. [Link]
ChangeExp [80] 2024.06 LLaVA-1.5 LLaVA-1.5 Studied the ability of large vision-language models (LVLM) to explain temporal changes in satellite images; proposed three prompting methods.
MAF-Net [81] 2024.07 ResNet-101 Tr-Dec Resnet extracts multi-scale bi-temporal features, and each scale’s bi-temporal features are processed and fused by Transformer.
SFEN [82] 2024.07 WideResNet Tr-Dec Proposed a scale feature enhancement network.
MfrNet [83] 2024.09 ResNet-18 Tr-Dec Introduced Joint Attention and Dense Feature Fusion Module (JADF) for refining features and reducing noise
SEIFNet [84] 2024.09 ResNet-101 Tr-Dec Proposed a network with cross-time interaction and symmetric difference learning to model differences from coarse to fine representations. [Link]
MV-CC [85] 2024.10 InternVideo2 Tr-Dec Proposed a mask-enhanced video model for change captioning to reduce the complex manual design of features extraction and fusion. [Link]
Chareption [86] 2024.10 CLIP ViT-L/14 LLaMA-7B Adjusted VIT and LLM through adapters to generate accurate and contextually relevant descriptions of changes.
MADiffCC [87] 2024.11 Diffusion Tr-Dec Use a diffusion feature extractor to capture the multi-level and bi-temporal features. Proposed a gated cross-attention guided captioning decoder.
CCExpert [88] 2024.11 ViT Qwen-2 Utilized the MLLM and designed a Difference-aware Integration Module to capture fine-grained differences and constructed a CC-Foundation dataset, comprising 200,000 pairs of images and 1.2 million captions. [Link]
Change3D [89] 2025.03 X3D-L(video) Tr-Dec Change3D reconstructs the bi-temporal change detection and description task from a video modeling perspective, achieving joint spatiotemporal feature learning by introducing learnable perception frames. [Link]
RDD+ACR [90] 2025.04 ResNet-101 Tr-Dec Proposed the RDD regional feature distillation framework and the ACR alignment mechanism, which enhanced the model’s ability to capture subtle regional changes through a ”concrete-to-abstract” two-stage learning strategy and alignment optimization between linguistic and visual differences.

### v3

# Remote Sensing SpatioTemporal Vision-Language Models: A Comprehensive Survey

@@ Abstract
The interpretation of multi-temporal remote sensing imagery is critical for monitoring Earth’s dynamic processes. However, previous change detection methods, which produce binary or semantic masks, fall short of providing human-readable insights into changes. Recent advances in Vision-Language Models (VLMs) have opened a new frontier by fusing visual and linguistic modalities, enabling spatiotemporal vision-language understanding: models that not only capture spatial and temporal dependencies to recognize changes but also provide a richer interactive semantic analysis of temporal images (e.g., generate descriptive captions and answer natural-language queries). In this survey, we present the first comprehensive review of RS-STVLMs. The survey covers the evolution of models from early task-specific models to recent general foundation models that leverage powerful large language models. We discuss progress in representative tasks, such as change captioning, change question answering, and change grounding. Moreover, we systematically dissect the fundamental components and key technologies underlying these models, and review the datasets and evaluation metrics that have driven the field. By synthesizing task-level insights with a deep dive into shared architectural patterns, we aim to illuminate current achievements and chart promising directions for future research in spatiotemporal vision-language understanding for remote sensing. We will keep tracing related works at https://github.com/Chen-Yang-Liu/Awesome-RS-SpatioTemporal-VLMs
@@
Remote Sensing, Spatiotemporal Understanding, Vision-Language Model, Foundation Model, and Large Language Model.
@@ I Introduction
Early efforts in remote sensing image interpretation primarily concentrated on single-image analysis tasks, such as land classification [9, 10, 11], object detection [12, 13], and semantic segmentation [14, 15, 16]. However, single-temporal imagery reflects only the spatial surface conditions at a specific moment and fails to capture dynamic change processes like urban expansion, deforestation, or disaster aftermath. As sensing platforms and revisit frequencies have improved [17, 18, 19, 20, 21], multi-temporal imagery has become available, opening the door to remote sensing spatiotemporal image interpretation: models must understand spatial patterns (e.g., object class, spatial context) and capture temporal evolution (e.g., appearance, disappearance, growth) across time.
@@
Early research on the spatiotemporal image interpretation focused on binary change detection, which produces pixel-level masks to localize changed object regions based on temporal image sequences, such as vegetation cover changes or the emergence of buildings [22, 23]. While effective at localizing where change occurs, these approaches lack semantic insight into what has changed. Semantic change detection extended this paradigm by assigning change object labels to changed areas [24, 25, 26], yet remained confined to visual-level assessments. They struggle to achieve a higher-level semantic understanding of changes and interactive queries in a human-readable manner, such as the evolving state of objects, color of changed objects and their position relationship [27, 28, 29].
@@
In pursuit of richer and more flexible semantic understanding, some researchers have turned to Remote Sensing SpatioTemporal Vision-Language Models (RS-STVLMs), which integrate multi-temporal imagery with natural language processing for dynamic scene understanding. RS-STVLMs covers multiple tasks such as change captioning [28, 27] and change question answering [30, 31], thereby enriching the toolkit for remote sensing spatiotemporal image interpretation. Language, as a vehicle for human communication and knowledge [32], enhances the higher-level understanding ability of RS-STVLMs. By combining language with temporal images, RS-STVLMs can not only identify spatiotemporal changes but also generate detailed descriptions and answers, facilitating multimodal interactions that extend beyond mere visual change detection.
@@
Fig. 1 illustrates the timeline of some representative RS-STVLMs along with their publication dates, highlighting an active research trend since 2021. Despite the promising advancements, existing research remains fragmented, typically concentrating on isolated, task-specific methods. This makes it challenging for researchers to obtain a holistic view of the field’s progress and emerging trends. There is a growing need for a holistic and systematic survey of RS-STVLMs.
From the perspective of vision-language models (VLMs), previous studies have explored various VLMs, such as RSGPT [33], GeoChat [34], H2RSVLM [35], LHRS-Bot [36], and EarthGPT [37]. Several surveys [38, 39, 40] have reviewed the development of VLMs within the remote sensing community. While these studies have made significant contributions, they primarily focus on static, single-temporal image tasks such as image captioning [41, 42, 43], visual question answering (VQA) [44, 45, 46], text-to-image retrieval [47, 48, 49, 50], and visual grounding [51, 52, 53]. In contrast, our survey focuses specifically on RS-STVLMs, a novel and emerging area that integrates multi-temporal imagery with natural language processing for dynamic scene understanding. Unlike conventional VLMs that focus on single-temporal images, RS-STVLMs are tasked with capturing both spatial changes and their temporal evolution over time. This introduces unique challenges, such as spatiotemporal modeling and cross-time semantic alignment, which are not addressed by conventional VLMs.
@@
Figure 1: Timeline of RS-STVLMs, illustrating the field’s rapid ascent. We outline the evolution of representative tasks and show advanced progress of how Large Language Models (LLMs) are reshaping the RS-STVLM landscape. For additional resources and daily updates, visit our GitHub page at https://github.com/Chen-Yang-Liu/Awesome-RS-SpatioTemporal-VLMs
In this survey, we present the first comprehensive review of RS-STVLMs. Fig. 2 shows the structure of this survey. The survey covers the evolution of models from early task-specific models to recent general foundation models that leverage powerful large language models (LLMs). We discuss progress in representative tasks, such as change captioning, change question answering, and change grounding. Moreover, we systematically dissect the fundamental components and key technologies underlying these models, and review the datasets and evaluation metrics that have driven the field. By synthesizing task-level insights with a deep dive into shared architectural patterns, we aim to illuminate current achievements and chart promising directions for future research in spatiotemporal vision-language understanding for remote sensing.
@@
Figure 3: Three key concepts in the remote sensing spatiotemporal image interpretation: (a) Binary change detection, locating areas of change without providing deeper contextual information. (b) Semantic change detection, offering insights into the types of changed objects. (c) SpatioTemporal Vision-Language Understanding, incorporating higher-level flexible language rather than being limited to visual-level interpretation.
II Evolution of SpatioTemporal Vision-Language Models
II-A From Visual Change Detection to SpatioTemporal Vision-Language Understanding
@@ II-A1 Remote Sensing Change Detection
Remote sensing change detection (CD) is a fundamental spatiotemporal image interpretation task aimed at identifying differences between images acquired at different time points [54, 55, 56]. The evolution of CD methods can be viewed from both a task and a technical perspective. In the early stages, the focus was on binary CD, locating areas of change at the pixel level without providing semantic insights into the specific types or attributes of the changes. As the field progressed, semantic CD emerged. It extends binary methods by labeling each changed region with a specific land-cover category, thereby providing a higher level of semantic interpretation.
On the perspective of technological evolution, early CD approaches relied on traditional techniques, including algebra-based methods such as Change Vector Analysis (CVA)[57, 58], transformation-based methods like Principal Component Analysis (PCA)[59, 60] and Multivariate Alteration Detection (MAD)[61, 62], as well as classification-based approaches such as post-classification comparisons[63, 64]. These conventional methods laid the groundwork for CD but were limited by their reliance on hand-crafted features and simple statistical models.
@@
CNN-based Methods have been the cornerstone of deep learning approaches for CD in remote sensing [65, 66, 67]. These models leverage hierarchical convolutional operations to extract spatial change features from bi-temporal images, making them effective at capturing local textures and patterns for refined pixel-level mask prediction. Daudt et al. [65] first introduced FCNs into the CD task and presented Siamese architectures to adapt bi-temporal images. To improve CD accuracy, many CNN-based works have been developed, such as proposing multi-scale feature fusion strategies [68, 69, 70], and designing attention modules [71, 72, 73].
@@
Transformer-based Methods represent a significant leap forward in CD [74, 75, 76, 77, 78, 79]. Different from CNNs with limited receptive fields, through the self-attention mechanism, the Transformer captures long-range dependencies and contextual relationships across the entire image. As a pioneering approach, Chen et al. [74] represent the multi-temporal images as a few semantic tokens and use a Transformer to refine the visual features, revealing the change of interest. Subsequently, many methods have emerged, such as using pure Transformer [80, 81], combining transformers and CNNs to extract rich global context and local detailed information [82, 83], and lightweight Transformer [77, 84, 85]
@@
Mamba-based methods represent a recent advantage in CD [86, 87, 88, 89, 90, 91]. Unlike Transformer, which typically exhibits quadratic complexity to the input sequence length, Mamba achieves linear complexity while maintaining a global receptive field. ChangeMamba [87] first introduced Mamba architecture for the CD task and proposed three spatial-temporal relationship modeling mechanisms. CDMamaba [88] integrates local clues in the Mamba block and dynamically performs the interaction of bi-temporal features guided by global-local features to achieve fine-grained CD. Zhao et al. [86] presented an Omnidirectional Selective Scan Module to selectively scan high-resolution images in multiple directions.
@@
In summary, the evolution from binary CD to semantic CD reflects a growing demand for a more comprehensive understanding of spatiotemporal dynamics in remote sensing imagery. Technically, the field has transitioned from traditional methods to deep learning approaches, progressing from CNN-based methods to Transformer-based models, and now to emerging frameworks like Mamba. These advancements have significantly improved the capacity to detect not only where changes occur, but also what types of changes occur, thereby laying the foundation for subsequent multimodal developments in RS-STVLMs.
II-A2 Remote Sensing SpatioTemporal Vision-Language Tasks
@@
Building on the advances in remote sensing CD, recent research has expanded towards integrating both visual and linguistic modalities, forming a set of comprehensive tasks collectively termed Temporal Vision-Language Understanding tasks. These tasks not only involve detecting and localizing changes in temporal images but also involve generating natural language descriptions, and interactive question answering. This evolution reflects the community’s pursuit of a deeper semantic understanding of dynamic changes in remote sensing spatiotemporal image analysis. In this section, we outline the primary tasks in RS-STVLMs, providing formal definitions.
@@
Formally, let $I=\{I_{t_{1}},I_{t_{2}},...,I_{t_{n}}\}$ represent a sequence of remote sensing images acquired at different time steps, and let $Q$ denote a textual query, prompt, or instruction. The goal of spatiotemporal vision-language models is to infer a task-specific output, which varies across different tasks.
@@
Others: A few pioneering works have explored novel spatiotemporal vision-language tasks. Text-driven Temporal Image Generation aims to generate a sequence of remote sensing images that reflects dynamic changes described in a user-provided textual prompt, such as [92, 93]. Irvin et al. [94] introduce Temporal Referring Expression task, which requires the model to identify when (in which image) a specific change described by the user occurs.
These tasks underscore the transition from early change detection methods that focus solely on visual cues to a multi-modal approach that leverages language for deep semantic understanding. By incorporating natural language, RS-STVLMs can provide richer, more human-centered insights into dynamic Earth observation.
II-B Remote Sensing SpatioTemporal Vision-Language Models
In this section, we introduce the methodologies underlying each of the above spatiotemporal vision-language understanding tasks, highlighting key architectures.
Figure 4: The general framework for some representative remote sensing spatiotemporal vision-language tasks: (a) Remote Sensing Change Captioning, (b) Multi-task learning of Change Detection and Captioning, (c) Remote Sensing Change Question Answering, (d) Remote Sensing Text-driven Temporal Images Retrieval, and (e) Remote Sensing Change Grounding.
@@ II-B Remote Sensing SpatioTemporal Vision-Language Models
CNN-RNN [95] 2021.10 VGG-16 RNN First explored the RS change captioning task on a private dataset.
CC-RNN/SVM [96] 2022.08 VGG-16 RNN, SVM Released two small-scale datasets and proposed two bi-temporal feature fusion strategies.
RSICCformer [28] 2022.11 ResNet-101 Tr-Dec Released the large-scale LEVIR-CC dataset and first proposed a dual-branch Transformer model for change captioning. [Link]
PSNet [97] 2023.07 ViT-B/32 Tr-Dec Designed difference-aware layers and scale-aware reinforcement modules for multi-scale feature extraction and progressive utilization. [Link]
PromptCC [98] 2023.10 ViT-B/32 GPT-2 Decoupled change captioning into binary change classification and fine-grained change perception, and proposed a multi-prompt learning strategy to guide LLMs in generating plausible captions. [Link]
Chg2Cap [27] 2023.11 ResNet-101 Tr-Dec Proposed an attentional encoder to precisely localize the changes between bi-temporal images. [Link]
ICT-Net [99] 2023.11 ResNet-101 Tr-Dec Proposed an Interactive Change-Aware Encoder combined with ResNet for multi-scale feature extraction and change representation. [Link]
SITS-CC [100] 2024.03 ResNet-101 Tr-Dec Explored a new task of image time-series change captioning. [Link]
RSCaMa [101] 2024.05 ViT-B/32 Mamba, Tr-Dec, GPT-2 Introduced Mamba to the task, developed an SSM-based spatiotemporal joint modelling strategy, and compared three different language decoders [Link]
SparseFocus [85] 2024.05 ResNet-101 Tr-Dec Proposed Sparse Focus Transformer to enhance attention to changing regions while reducing computational complexity. [Link]
SEN [102] 2024.05 ResNet with 6-channel Tr-Dec Introduced a Single-stream Extractor with contrastive pre-training, achieving lower computational cost compared to dual-stream extractors. [Link]
Diffusion-RSCC [103] 2024.05 ResNet-101 Diffusion Applied diffusion models to learn cross-modal distributions between image pairs and change captions. [Link]
@@
ChangeRetCap [105] 2024.06 ResNet-101 Tr-Dec Proposed a foundation model addressing both bi-temporal captioning and bi-temporal text-guided temporal image retrieval. [Link]
ChangeExp [106] 2024.06 LLaVA-1.5 LLaVA-1.5 Investigated the capability of large vision-language models in explaining temporal changes and explored three prompting strategies.
MAF-Net [107] 2024.07 ResNet-101 Tr-Dec Employed ResNet for extracting multi-scale bi-temporal features, which were then fused at each scale using a Transformer-based module with gated attention for language decoding.
SFEN [108] 2024.07 WideResNet Tr-Dec Proposed a Scale Feature Enhancement Network to improve multi-scale representation learning and filter redundant bitemporal features.
MfrNet [109] 2024.09 ResNet-18 Tr-Dec Introduced a Joint Attention and Dense Feature Fusion Module (JADF) to refine features and suppress noise.
SEIFNet [110] 2024.09 ResNet-101 Tr-Dec Designed a network incorporating cross-temporal interaction and symmetric difference learning to capture hierarchical change representations (from coarse to fine). [Link]
MV-CC [111] 2024.10 InternVideo2 Tr-Dec Proposed a mask-enhanced video model to simplify the feature extraction and fusion process in change captioning. [Link]
Chareption [112] 2024.10 CLIP ViT-L/14 LLaMA-7B Leveraged adapter-tuned ViT and LLM to generate accurate and context-aware change descriptions.
MADiffCC [113] 2024.11 Diffusion Tr-Dec Employed a diffusion-based extractor to model bi-temporal, multi-level representations and introduced a gated cross-attention captioning decoder.
CCExpert [114] 2024.11 ViT Qwen-2 Leveraged a multi-modal large language model and designed a Difference-aware Integration Module to detect fine-grained changes. [Link]
Change3D [115] 2025.03 X3D-L(video) Tr-Dec Reconstructed the change detection and captioning task from a video modeling perspective via learnable perception frames for joint spatiotemporal representation learning. [Link]
RDD+ACR [116] 2025.04 ResNet-101 Tr-Dec Proposed region-aware difference distilling network with attribute-guided contrastive regularization to enhance fine-grained alignment between linguistic and visual modalities.